OpenStudy (m):

change from rectangular to spherical coordinates (0, √3, 1)

OpenStudy (amistre64):

spherical; ... as that the one thats relate to polar? or the one with phi in it?

OpenStudy (m):

phi theta

OpenStudy (amistre64):

one uses r the other uses angles

OpenStudy (m):

i think distance formula for p

OpenStudy (amistre64):

ditance = 2 :)

OpenStudy (amistre64):

2, gonna have to look it up to refresh lol

OpenStudy (m):

ok i get it. you just need to plug in to formulas after that

OpenStudy (amistre64):

r, theta, phi; r = distance from origin; theta = tan-1(y/x); phi = cos-1(z/r)

OpenStudy (amistre64):

(2,pi/2,pi/4) right?

OpenStudy (m):

pi/3 for last one

OpenStudy (amistre64):

z=1 tho' 1/2

OpenStudy (amistre64):

your right; i was thinking 1/sqrt(2) lol

OpenStudy (m):

yeah hehe